40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14466     Entry time: Tue Feb 19 22:52:17 2019
Author: gautam 
Type: Update 
Category: ASS 
Subject: Y arm clipping doubtful 

In an earlier elog, I had claimed that the suspected clipping of the cavity axis in the Y arm was not solved even after shifting the heater. I now think that it is extremely unlikely that there is still clipping due to the heater. Nevertheless, the ASS system is not working well. Some notes:

  1. The heater has been shifted nearly 1-inch relative to the cavity axis compared to its old position - see Attachment #1 which compares the overhead shot of the suspension cage before and after the Jan 2019 vent.
  2. On Sunday, I was able to recover TRY ~ 1.0 (but not as high as I was able to get by intentionally setting a yaw offset to the ASS) by hand alignment with the spot on ETMY much closer to the center of the optic, judging by the camera. There are offsets on the dither alignment error signals which depend on the dither frequency, so the A2L signals are not good judges of how well centered we are on the optic.
  3. By calculating the power lost by clipping a Gaussian beam cross-section with a rectangular block from one side (an admittedly naive model of clipping), I find that we'd have to be within 15 mm of the line connecting the centers of ITMY and ETMY to even see ~10 ppm loss, see Attachment #2. So it is hard to believe that this is still a problem. Also, see  Attachment #3 which compares side-by-side the view of ETMY as seen through the EY optical table viewport before and after the Jan 2019 vent.

We have to systematically re-commission the ASS system to get to the bottom of this.

Attachment 1: overheadComparison.pdf  28.072 MB  | Hide | Hide all
overheadComparison.pdf
Attachment 2: clipping.pdf  14 kB  | Hide | Hide all
clipping.pdf
Attachment 3: rearComparison.pdf  9.903 MB  | Hide | Hide all
rearComparison.pdf
ELOG V3.1.3-